# CodePMP: Scalable Preference Model Pretraining for Large Language Model Reasoning

### v1

@@ 1 Introduction
Large language models (LLMs) have made remarkable progress in natural language understanding and generation, benefiting from scalable pretraining and finetuning techniques like supervised finetuning (SFT) (Wang et al., 2022; 2023a) and Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022a; Lightman et al., 2023b; Bai et al., 2022b; Gulcehre et al., 2023; Schulman et al., 2017; Rafailov et al., 2024). However, enhancing LLMs’ reasoning abilities, particularly in complex logical and mathematical tasks, remains a significant challenge (Wang et al., 2023b; Zhang et al., 2024b). Although RLHF is effective, it relies heavily on high-quality preference data, which is costly and labor-intensive to annotate (Cobbe et al., 2021b; Zheng et al., 2024). This limitation impedes the scalability of reward model (RM) finetuning, which is essential for guiding LLMs toward optimal outputs.
@@
To alleviate this issue, prior works like Anthropic’s Preference Model Pretraining (PMP) (Askell et al., 2021) have proposed improving reward modeling data efficiency by pretraining preference models on large-scale preference data from public sources like Reddit and Wikipedia, followed by an efficient finetuning on limited high-quality human-annotated data. However, this approach is less effective for reasoning tasks due to the scarcity of reasoning preference pairs available online. Compared to other tasks, manually annotating preference data for reasoning is inherently more challenging and difficult to scale (Zhang et al., 2024b; Zhou et al., 2023), highlighting the urgent need for a scalable PMP approach for reasoning tasks.
Figure 2: An illustration of the CodePMP process. First, raw code collected from GitHub is cleaned and summarized into code prompts (descriptions). Then, for each code prompt, a weak CodeLLM generates a rejected response, while a stronger CodeLLM produces a chosen response. Finally, these <chosen, rejected> pairs, accumulated in the millions, form the pretraining dataset for the preference model. This pretraining process improves not only sample efficiency but also the performance for downstream reasoning reward model finetuning.
@@
In this paper, we propose CodePMP, a scalable preference model pretraining pipeline that enhances LLM reasoning abilities using synthesized preference pairs derived from high-quality, publicly available source code. Code, with its inherently logical and structured nature, provides rich data suitable for reasoning tasks. Recent works (Zhang et al., 2024b; Aryabumi et al., 2024) also show a strong correlation between code training and reasoning improvements in LLMs. By leveraging the huge amount and diverse coverage of source code available on platforms like GitHub, CodePMP offers a scalable solution for pretraining preference models, thereby improving RM finetuning efficiency and enhancing LLMs reasoning performance.
Specifically, CodePMP generates preference pairs by synthesizing chosen and rejected code responses for a given code-related prompt or description using CodeLLMs. A strong CodeLLM produces higher-quality (chosen) responses, while a weaker model generates suboptimal or even low-quality (rejected) responses. These <chosen, rejected> pairs, accumulated in the millions, form a large-scale synthesized preference dataset. This dataset is then used to pretrain the preference model with pairwise ranking objectives (Cobbe et al., 2021b; Charniak & Johnson, 2005), providing an good initialization for further fine-tuning the reward models.
@@
We evaluate CodePMP on widely studied reasoning tasks, including mathematical reasoning tasks such as GSM8K (Cobbe et al., 2021b) and MATH (Hendrycks et al., 2021), as well as logical reasoning tasks like ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023). Our experiments demonstrate that CodePMP significantly improves RM fine-tuning accuracy and Best-of-N performance in reasoning tasks, outperforming direct RM fine-tuning, as highlighted in Figure 1. Moreover, additional experimental results reveals that RMs initialized with CodePMP exhibit greater robustness across different tasks. These results indicate that code-derived preference data provides a scalable, cost-effective solution for enhancing LLM reasoning capabilities while reducing reliance on extensive preference annotation, achieving more effective reward modeling for reasoning tasks.
@@ 2 Preliminaries
Language modeling (LM) is a fundamental task in natural language processing, aimed at modeling sequences of language. This is typically achieved through Causal Language Models (Causal LM), where the model is trained to maximize the likelihood of predicting the next word $w_{t}$ given the preceding words $w_{1},w_{2},\dots,w_{t-1}$. The training process minimizes the negative log-likelihood of the predicted word sequence:
$\mathcal{L}_{\text{LM}}=-\sum_{t=1}^{T}\log P(w_{t}|w_{1},w_{2},\dots,w_{t-1})$
This loss function $\mathcal{L}_{\text{LM}}$ encourages the model to capture the underlying patterns in the data. Transformer architectures (Vaswani, 2017) have become the standard for Causal LM due to their ability to handle long-range dependencies effectively.
@@
Reward modeling (RM) is crucial in reinforcement learning from human feedback (RLHF), providing a scalar reward signal that guides the learning process based on output quality. The reward model $R_{\theta}$ predicts the quality of an output $y$ given a context $x$ as $s=R_{\theta}(x,y)$, where $s$ is the scalar reward score. In preference modeling, RM predicts the relative quality of outputs by comparing pairs. A common method is the Pairwise Ranking Loss, where the model assigns higher scores to preferred (chosen) outputs:
$\mathcal{L}_{\text{RM}}=-\log\left(\sigma(s_{\text{chosen}}-s_{\text{rejected}})\right)$
@@
Best-of-N (BoN) sampling improves LLM reasoning (Cobbe et al., 2021b; Lightman et al., 2023b). In this approach, $N$ candidate solutions $\{y_{1},y_{2},\dots,y_{N}\}$ are generated by sampling from the LLM’s output distribution for a given problem. A reward model scores each candidate and selects the highest-scoring one as the final answer:
$\hat{y}=\arg\max_{y_{i}\in\{y_{1},y_{2},\dots,y_{N}\}}R_{\theta}(x,y_{i})$
@@ 3.1 Model Design
Code Preference Model Pretraining (CodePMP) is designed to enhance the sample efficiency of reward models, particularly for reasoning tasks where high-quality preference data is scarce. Traditionally, reward models are finetuned on small, curated datasets, which limits their effectiveness in complex tasks like mathematical reasoning or logical deduction. CodePMP mitigates this limitation by introducing a pretraining phase between basic language model pretraining and finetuning on domain-specific reasoning datasets. This phase leverages a large, diverse dataset of code-preference pairs, enabling the model to learn generalizable patterns and ranking strategies.
CodePMP training involves two key components: Reward Modeling (RM) and Language Modeling (LM). In RM, the model is trained on code-preference pairs, learning to assign higher scores to the chosen code through a pairwise ranking loss. In LM, only the chosen code is used for autoregressive training to maintain the model’s general capabilities. The overall loss is a combination of the RM and LM losses, ensuring the model enhances its ranking ability without sacrificing general language modeling performance: $\mathcal{L}_{\text{PMP}}=\mathcal{L}_{\text{rank}}+\mathcal{L}_{\text{LM}}$.
@@ 3.2 Data Construction
To enable scalable preference model pretraining, we construct a dataset sourced from GitHub, which includes a diverse range of repositories and associated metadata. The dataset consists of two primary components: Repository Data comprises over 1.3 billion code files from GitHub repositories, while GitHub Metadata includes information such as commit histories, discussions, pull requests, and issues.
The CodePMP dataset is constructed through a systematic process. First, raw source code is processed by a description summarizer, typically an instruction-tuned CodeLLM, to generate prompts that describe the functionality of the code.
These prompts are then used by two CodeLLMs of different capabilities to generate code snippets:
@@
This process yields pairs of code responses—one chosen and one rejected—which are used for preference modeling. This scalable approach significantly enhances pretraining efficiency, improving performance on downstream tasks.
The steps of the CodePMP methodology are outlined systematically in Algorithm 1.
Algorithm 1 Code Preference Model Pretraining (CodePMP)
1:Source code repository $S$, Strong CodeLLM $M_{\text{strong}}$, Weak CodeLLM $M_{\text{weak}}$
2:Pretrained Model
3:Input: Source code $S$
4:Summarize description $D$ using $M_{\text{strong}}$ on $S$
5:for each $D_{i}\in D$ do
6:     Generate Chosen Response using $M_{\text{strong}}$
7:     Generate Rejected Response using $M_{\text{weak}}$
8:end for
9:Calculate LM Loss $\mathcal{L}_{\text{LM}}$ on Response
10:Calculate RM Loss $\mathcal{L}_{\text{rank}}$ using Chosen Response and Rejected Response
11:Train PMP Model using $\mathcal{L}_{\text{PMP}}=\mathcal{L}_{\text{rank}}+\mathcal{L}_{\text{LM}}$
4 Experimental
In this section, we first outline the experimental setup, followed by the experimental results, highlighting that CodePMP is a highly scalable method.
@@ 4.1.1 CodePMP Settings
We generate code preference pairs following Algorithm 1, using the deepseek-coder-6.7b-instruct model as the strong CodeLLM to generate chosen responses and the deepseek-coder-1.3b-instruct model as the weak CodeLLM to generate rejected responses. The constructed CodePMP dataset includes 28 million files and 19 billion tokens. The diverse datasets provide sufficiently broad prompt coverage for preference model pretraining, which is conducive to the generalization of preference models in reasoning tasks. In addition, the average lengths of the chosen and rejected responses are similar, ensuring that response length does not bias the CodePMP learning process. Details are provided in Appendix C and Table 4.
@@
By default, we initialize the preference models with the publicly available Qwen models (Yang et al., 2024a), using different model sizes, specifically Qwen2-1.5B and Qwen2-7B. Detailed hyperparameters for CodePMP training are provided in Appendix B.
@@ 4.1.2 Reasoning Finetuning Settings
We validate CodePMP on reward models across two reasoning task types: mathematical and logical reasoning. The reward model is finetuned on corresponding preference datasets for each task. For mathematical reasoning, we use the MathShepherd-pair dataset, derived from MathShepherd (Wang et al., 2023b), and evaluate the model on a holdout test set to assess RM accuracy.
Similarly, for logical reasoning, we use the ReClor-pair and LogiQA2.0-pair datasets, derived from ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023), respectively. We train reward models on these datasets, with holdout test sets used to evaluate model accuracy. Dataset construction and finetuning hyperparameters are provided in Appendix D and  B.
@@ 4.1.3 Evaluation Settings
Following (Zhang et al., 2024a), we adopt two evaluation metrics:
RM Accuracy
This metric measures the accuracy of the reward model in distinguishing chosen from rejected solutions on the holdout test sets. It provides insight into the model’s ability to classify individual sequences.
Best-of-N (BoN) Accuracy
This metric evaluates the proportion of correct solutions selected by the finetuned RM from $N$ candidate responses. It assesses the model’s group-wise ranking performance, focusing on its ability to select the correct answer from a set of candidates. We use MetaMath-Mistral-7B (Yu et al., 2023) as the generator for this evaluation.
For mathematical reasoning, we use the GSM8K (Cobbe et al., 2021b) and MATH (Hendrycks et al., 2021) test sets. For logical reasoning, we evaluate on the ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023) test sets. Further details can be found in Appendix D.
Note that logical reasoning questions typically involve a paragraph followed by statements to be judged true or false, making Best-of-N evaluation challenging. Therefore, we use multiple-choice accuracy, where the reward model ranks four manually annotated options and selects the best one. This metric is equivalent to Best-of-4, and thus, for logical reasoning tasks, multiple-choice accuracy and Best-of-N are used interchangeably.
@@ 4.2.1 RM Accuracy Results
Table 1: Comparison of RM accuracies: reward models finetuned with CodePMP initialization achieve higher accuracies on the reasoning holdout test sets, demonstrating an improved ability to distinguish chosen responses from rejected ones.
Model PMP MathShepherd-pair Reclor-pair LogiQA2.0-pair
1.5B 0.7226 0.758 0.7538
0.8186 0.794 0.7774
7B 0.8777 0.862 0.8263
0.9274 0.874 0.8441
@@
Figure 3: Comparison of Best-of-N accuracies: reward models finetuned with CodePMP initialization consistently perform better, with improvements remaining robust as N increases, demonstrating CodePMP’s effectiveness in improving group-wise ranking capabilities.
(a) GSM8K / 1.5B
(b) MATH / 1.5B
(c) Reclor / 1.5B
(d) LogiQA2.0 / 1.5B
(e) GSM8K / 7B
(f) MATH / 7B
(g) Reclor / 7B
(h) LogiQA2.0 / 7B
Figure 4: Comparison of sample efficiency in RM finetuning: reward models finetuned with CodePMP initialization consistently achieve substantially higher Best-of-N accuracy when finetuning on the same amount of samples, demonstrating superior sample efficiency. Note that the horizontal axis grows exponentially with $\sqrt{2}$. The green lines represent the settings with CodePMP, while the blue lines represent the settings without CodePMP.
@@ 4.2.2 BoN Accuracy Results
We evaluate BoN accuracy across reasoning tasks with and without CodePMP initialization. As shown in Figure 8(a) and Table 8(b), RM finetuned with CodePMP initialization consistently achieves higher BoN accuracy across both mathematical and logical reasoning tasks for 1.5B and 7B models. This highlights CodePMP’s effectiveness in improving RM’s group-wise ranking performance.
Across different values of N, RM models initialized with CodePMP maintain their lead, showing robust improvement even as N increases to 256. In contrast, RM without CodePMP shows a sharp decline in accuracy as N increases, underscoring the stability CodePMP provides, likely due to the diverse code-preference pairs used during training.
For logical reasoning, the performance gap between CodePMP and non-PMP models is smaller, as $N$ is limited to 4, while in mathematical reasoning, $N$ reaches 256. This suggests that increasing $N$ further in logical reasoning could amplify the advantages of CodePMP in future evaluations.
4.2.3 RM Sample Efficiency Comparison
One key advantage of CodePMP is its ability to improve the sample efficiency of RM finetuning. To assess this, we conduct experiments with progressively larger sample sizes for RM finetuning. As indicated by (Kaplan et al., 2020), optimal results are achieved when the learning rate scheduler completes its decay at the end of training. Therefore, rather than evaluating intermediate checkpoints, we retrain models with varying sample sizes for optimal results. Figure 4 and  10 (Appendix E) show that as the sample size increases, RMs with CodePMP initialization consistently outperforms others in both BoN and RM accuracy. Notably, RMs finetuned with CodePMP initialization using just 0.5k samples surpasses RMs finetuned without CodePMP initialization using 40k samples on mathematical tasks, demonstrating CodePMP’s significant advantage in sample efficiency. However, as sample size increases, this advantage diminishes slightly, suggesting that with much larger datasets, CodePMP’s benefit may become less pronounced, but the cost of manual labeling remains a key consideration.
4.2.4 The Importance of Scalable PMP

### v2

@@ Abstract
CodePMP: Scalable Preference Model Pretraining for Large Language Model Reasoning
Huimu Yu1,2, Xing Wuthanks: Equal contribution. This work is jointly completed by Huimu Yu under the guidance of Xing Wu.1,2,3 🖂, Haotian Xu3, Debing Zhang3, Songlin Hu1,2 🖂 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences 3Xiaohongshu Inc {yuhunmu,wuxing,husonglin}@iie.ac.cn, {xuhaotian,dengyang}@xiaohongshu.com
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress in natural language understanding and generation, driven by advancements in scalable pretraining and finetuning techniques, including supervised finetuning (SFT) (Wang et al., 2022, 2023a) and Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022a; Lightman et al., 2023; Bai et al., 2022b; Gulcehre et al., 2023; Schulman et al., 2017; Rafailov et al., 2024). Despite these advances, enhancing LLMs’ reasoning capabilities, particularly for complex logical and mathematical tasks, remains a significant challenge (Wang et al., 2023b; Zhang et al., 2024b). While RLHF has proven effective for improving model performance, its efficacy is constrained by the availability of high-quality preference data, which is expensive and labor-intensive to collect (Cobbe et al., 2021; Zheng et al., 2024). This limitation impedes the scalability of reward model (RM) finetuning, which is instrumental in guiding LLMs toward optimal outputs.
@@
To alleviate this issue, prior works like Anthropic’s Preference Model Pretraining (PMP) (Askell et al., 2021) have proposed improving reward modeling data efficiency by pretraining preference models on large-scale preference data from public sources like Reddit and Wikipedia, followed by an efficient finetuning on limited high-quality human-annotated data. However, this approach is less effective for reasoning tasks due to the scarcity of reasoning preference pairs available online. Compared to other tasks, manually annotating preference data for reasoning is inherently more challenging to scale (Zhang et al., 2024b; Zhou et al., 2023), highlighting the urgent need for a scalable PMP approach for reasoning tasks.
[…]
@@
In this paper, we propose CodePMP, a scalable preference model pretraining pipeline that enhances LLM reasoning abilities using synthesized preference pairs derived from high-quality, publicly available source code. Code, with its inherently logical and structured nature, provides rich data suitable for reasoning tasks. Recent works (Zhang et al., 2024b; Aryabumi et al., 2024) also show a strong correlation between code training and reasoning improvements in LLMs. By leveraging the huge amount and diverse coverage of source code available on platforms like GitHub, CodePMP offers a scalable solution for pretraining preference models, thereby improving RM finetuning efficiency and enhancing LLMs’ reasoning performance.
[…]
@@
We evaluate CodePMP on widely studied reasoning tasks, including mathematical reasoning tasks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), as well as logical reasoning tasks like ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023). Our experiments show that CodePMP significantly improves RM finetuning accuracy and Best-of-N performance in reasoning tasks, outperforming direct RM finetuning, as highlighted in Figure 1. Moreover, additional results reveal that RMs initialized with CodePMP exhibit greater robustness across different tasks. These results indicate that code-derived preference data provides a scalable, cost-effective solution for enhancing LLM reasoning capabilities while reducing reliance on extensive preference annotation, achieving more effective reward modeling for reasoning tasks.
@@ 2 Preliminaries
Language modeling represents a fundamental task in natural language processing aimed at modeling sequential language data. This is typically implemented through Causal Language Models (Causal LM), which maximize the likelihood of predicting the next token $w_{t}$ given preceding tokens $w_{1},w_{2},\dots,w_{t-1}$. The training process minimizes the negative log-likelihood:
$\mathcal{L}_{\text{LM}}=-\sum_{t=1}^{T}\log P(w_{t}|w_{1},w_{2},\dots,w_{t-1})$ (1)
This loss function $\mathcal{L}_{\text{LM}}$ encourages the model to capture underlying patterns in the data. Transformer architectures (Vaswani, 2017) are the standard for Causal LM due to their ability to handle long-range dependencies effectively.
@@
Reward modeling (RM) is integral to reinforcement learning from human feedback (RLHF), providing scalar reward signals that guide learning based on output quality. The reward model $R_{\theta}$ predicts the quality of an output $y$ given a context $x$ as $s=R_{\theta}(x,y)$. In preference modeling, RMs predict relative quality by comparing output pairs. A standard approach employs the Pairwise Ranking Loss, which assigns higher scores to preferred (chosen) outputs:
$\mathcal{L}_{\text{RM}}=-\log\left(\sigma(s_{\text{chosen}}-s_{\text{rejected}})\right)$ (2)
@@
Best-of-N (BoN) sampling enhances LLM reasoning (Cobbe et al., 2021; Lightman et al., 2023) by generating $N$ candidate solutions $\{y_{1},y_{2},\dots,y_{N}\}$ for a given problem, then using a reward model to score and select the highest-scoring candidate:
$\hat{y}=\arg\max_{y_{i}\in\{y_{1},y_{2},\dots,y_{N}\}}R_{\theta}(x,y_{i})$ (3)
@@ 3.1 Model Design
Code Preference Model Pretraining (CodePMP) enhances the sample efficiency of reward models, particularly for reasoning tasks where high-quality preference data is scarce. Traditionally, reward models are finetuned on small, curated datasets, limiting their effectiveness in complex tasks like mathematical reasoning or logical deduction. CodePMP mitigates this limitation by introducing a pretraining phase between basic language model pretraining and finetuning on domain-specific reasoning datasets. This phase leverages a large, diverse dataset of code-preference pairs, enabling the model to learn generalizable patterns and ranking strategies.
CodePMP training involves two components: Reward Modeling (RM) and Language Modeling (LM). In RM, the model is trained on code-preference pairs, learning to assign higher scores to the chosen code through a pairwise ranking loss. In LM, only the chosen code is used for autoregressive training to maintain the model’s general capabilities. The overall loss combines the RM and LM losses, ensuring the model enhances its ranking ability without sacrificing general language modeling performance: $\mathcal{L}_{\text{PMP}}=\mathcal{L}_{\text{RM}}+\mathcal{L}_{\text{LM}}$.
@@ 3.2 Data Construction
To enable scalable preference model pretraining, we construct a dataset sourced from GitHub, containing over 1.3 billion code files from GitHub repositories. The CodePMP dataset is constructed through a systematic process. First, raw source code is processed by a description summarizer—typically an instruction-tuned CodeLLM—to generate prompts describing the code’s functionality. Two CodeLLMs with different capabilities then generate code snippets based on these prompts:
@@
This process yields pairs of code responses—one chosen and one rejected—which are used for preference modeling. This scalable approach significantly enhances pretraining efficiency, improving performance on downstream tasks. The steps of the CodePMP methodology are outlined systematically in Figure 2.
4 Experiments
In this section, we outline the experimental setup and then the experimental results, highlighting that CodePMP is a highly scalable method.
PMP
MathShepherd
-pair
Reclor
-pair
LogiQA2.0
-pair
Qwen2-1.5B
0.7226 0.758 0.7538
0.8186 0.794 0.7774
Qwen2-7B
0.8777 0.862 0.8263
0.9274 0.874 0.8441
Table 1: Reward model accuracy comparison: CodePMP-initialized models perform better on reasoning test sets, showing better discrimination ability.
@@ 4.1.1 CodePMP Settings
We generate code preference pairs by using the deepseek-coder-6.7b-instruct model as the strong CodeLLM to generate chosen responses and the deepseek-coder-1.3b-instruct model as the weak CodeLLM to generate rejected responses. The constructed CodePMP dataset includes 28 million files and 19 billion tokens. The diverse datasets provide sufficiently broad prompt coverage for preference model pretraining, which is conducive to the generalization of preference models in reasoning tasks. In addition, the average lengths of the chosen and rejected responses are similar, ensuring that response length does not bias the CodePMP learning process. Details are provided in Appendix.
@@
By default, we initialize the preference models with the publicly available Qwen models (Yang et al., 2024), using different model sizes, specifically Qwen2-1.5B and Qwen2-7B. Detailed hyperparameters for CodePMP training are provided in Appendix.
@@ 4.1.2 Reasoning Finetuning Settings
We evaluate CodePMP on mathematical and logical reasoning tasks using dedicated preference datasets. For mathematical reasoning, we finetune reward models on MathShepherd-pair dataset, derived from MathShepherd (Wang et al., 2023b), while logical reasoning models use ReClor-pair and LogiQA2.0-pair datasets, derived from ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023) respectively. Each model is finetuned on its corresponding training set and evaluated on its respective holdout test set for accuracy assessment. Implementation details for dataset construction and hyperparameters are provided in Appendix.
@@ 4.1.3 Evaluation Settings
Following (Zhang et al., 2024a), we evaluate using two metrics: (1) RM Accuracy measures the reward model’s ability to distinguish chosen from rejected solutions on holdout test sets, providing insight into the model’s ability to classify individual sequences; and (2) Best-of-N (BoN) Accuracy assesses the percentage of correct solutions selected by the RM from $N$ candidate responses, evaluating the model’s group-wise ranking performance and ability to identify the best answer from multiple candidates. We use MetaMath-Mistral-7B (Yu et al., 2023) as the generator for BoN evaluation.
We evaluate on GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) for mathematical reasoning, and ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023) for logical reasoning. For logical reasoning tasks, we use multiple-choice accuracy (equivalent to Best-of-4) where the RM ranks four manually annotated options, as logical reasoning questions typically consist of paragraphs followed by statements to be judged, making standard BoN evaluation challenging.
@@ 4.2.1 RM Accuracy Results
Figure 3: Best-of-N accuracy comparison: CodePMP-initialized models outperform baselines across various N values, showing superior ranking capabilities.
@@ 4.2.2 BoN Accuracy Results
Evaluations across reasoning tasks demonstrate that CodePMP-initialized RMs consistently achieve higher BoN accuracy on both mathematical and logical reasoning tasks for all model sizes (Figure 3). CodePMP models maintain performance advantages even as N increases to 256, while non-CodePMP models exhibit significant accuracy degradation at higher N values, highlighting CodePMP’s stability.
This aligns with research on BoN sampling (Chow et al., 2024) that identifies an inflection point where performance typically deteriorates beyond certain N thresholds due to increased base policy stochasticity and verifier misalignment. CodePMP-initialized models demonstrate greater stability at higher N values, suggesting improved alignment with true reward signals and enhanced robustness to noise amplification inherent in large-N sampling.
For logical reasoning, the performance gap appears smaller as testing was limited to N=4, while mathematical reasoning extended to N=256, suggesting potential for amplified advantages in logical reasoning with increased N values.
4.2.3 Sample Efficiency Analysis
To assess CodePMP’s impact on sample efficiency, we evaluated models with varying finetuning dataset sizes following best practices (Kaplan et al., 2020). Figure 4 shows that CodePMP-initialized models consistently outperform baselines across all dataset sizes, with CodePMP achieving with just 0.5k samples what baseline models require 40k samples to match—an 80× efficiency improvement. This advantage, while diminishing with larger datasets, significantly reduces annotation costs for developing effective reward models.
4.2.4 Scalability Analysis